# Communicate Less, Synthesize the Rest: Latency-aware Intent-based Generative Semantic Multicasting with Diffusion Models

### v2

@@ I Introduction
Recently, the Generative AI (GenAI) models have shown great success in developing efficient low bitrate Generative Semantic Communication (Gen Semcom) systems [2, 3, 4, 5]. Generative models are capable of learning the general distribution of natural signals during training and synthesizing new samples with high perceptual quality at inference time. The generation process can be guided by multi-modal prompts and conditioning signals to produce high quality outputs with a desired semantic content. The emerging Generative Foundation Models and Multimodal Large Language Models (MLLMs), e.g. Sora, Lumiere, and DALL.E, etc., provide ample opportunities to develop efficient generative Semcom frameworks [6, 7, 8, 9, 10]. As these models possess a vast general knowledge captured via intensive pre-training on huge amount of data, they alleviate the need for a shared Knowledge Base (KB) between the semantic transmitter and receiver, thereby reducing the corresponding knowledge sharing overheads. This vast general knowledge also makes the generative Semcom framework applicable to various datasets and tasks thereby achieving universality. The GenAI models are trained to maximize the perceptual quality of the synthesized signal, and the theoretical limits of generative Semcom are governed by the rate-distortion-perception theory [11, 12]. This theory explores the threefold trade-off between rate, distortion, and the perceptual quality of the synthesized signal. While the rate-distortion-perception function is analytically derived only for a few source distributions, it is generally estimated empirically for natural signals, e.g. images, audio/video, point cloud, etc., by training deep source encoders with a perceptual loss function, e.g. Wasserstein distance, on large datasets. In generative Semcom, the transmitter extracts the intended semantics, e.g. in form of textual prompts [13, 14, 15], compressed embeddings [16, 17], semantic map [18, 19], edge map [2, 7], etc., which are then transmitted over the channel. The receiver uses these semantics to guide a generative model, synthesizing a signal that is semantically consistent and highly realistic. Despite the above works, multiuser generative semantic communications, specifically in the multicasting setup, and its scalability to increased number of users, considering the communication/computation latency remains less studied.
[20] [21] [22] [23] This Work
@@
In this paper, we develop an scalable latency-aware generative semantic multicasting framework with adaptive intent-based resource allocation utilizing pre-trained diffusion models. We claim three key benefits for the proposed framework: Firstly, generative multicasting in our Semcom framework improves utilization of the wireless and network resources compared with non-generative and intent-unaware multicasting benchmarks. This also reduces the total latency while maintaining the reconstruction/synthesis distortion/perception quality. Secondly, the adoption of pre-trained models allows a source-channel separation-based Semcom architecture, thereby alleviating the need for end-to-end joint training of the transmitter and receiver, which is required in many existing Semcom frameworks [24, 25, 26, 27]. Such a separation-based architecture offers improved adaptability to varying channel conditions, improved scalability to increase the number of users, and better compatibility with the existing design of wireless communication networks. Finally, diffusion models enable synthesizing photorealistic outputs with better diversity, realism, and fine-grained details. TABLE I provides an overview of the major differences between this work and the key existing works on semantic multicasting/broadcasting [20, 21, 22, 23]. To the best of our knowledge, this work is the first of its kind to design joint communication/computation latency-aware intent-based generative semantic multicasting with pre-trained diffusion models, based on source-channel coding separation to achieve scalable and adaptable semantic multicasting compatible with the existing wireless networks, and derive the corresponding rate-distortion/perception tradeoffs. The contributions of this work are thereby three-fold:
@@ II-A1 Multi-class Semantic Decomposition with DDRNets
At the transmitter, a semantic segmentation model is used to generate a semantic map $\bf M$} from the source signal as depicted in Fig. 1 for sample image signals. To do this, any existing semantic segmentation model, e.g. [28, 29, 30], can be utilized, and the choice of the appropriate model depends on the required accuracy and affordable computational complexity. To minimize the end-to-end latency, we use a deep dual-resolution network (DDRNet)[29]-based architecture. DDRNets are efficient architectures that were proposed to achieve real-time performance for semantic segmentation of the street scenes. DDRNets are composed of two deep convolutional branches, and a deep information extractor called Deep aggregation pyramid pooling module (DAPPM) to enlarge effective receptive fields and fuse multi-scale context based on low-resolution feature maps as shown in Fig. 2. RB and RBB represent sequential and single residual bottleneck blocks. Information is extracted from low-resolution feature maps e.g., $1/8$ denotes high-resolution branch create feature maps whose resolution is 1/8 of the input image resolution. The black solid lines demonstrate the data processing path (including up-sampling and down-sampling) and black dashed lines denote information paths without data processing. The architecture is trained with the cross entropy loss to classify each image pixel into a semantic class, thereby generating the semantic map $\bf M$}. The semantic map is then converted to multiple one-hot binary masks ${\bf M}_{l}$. With this representation, we have ${\bf X}_{l}={\bf M}_{l}\odot{\bf X}$. The intended mask of user $k$ is then ${\bf M}^{(k)}_{I}=\{\bigcup_{l}{\bf M}_{l}|[{\bf I}]_{kl}=1\}$ and we have ${\bf X}^{(k)}_{I}={\bf M}^{(k)}_{I}\odot{\bf X}$. Similarly, the non-intended mask of user $k$ is then ${\bf M}^{(k)}_{N}=\{\bigcup_{l}{\bf M}_{l}|[{\bf I}]_{kl}=0\}$ and we have ${\bf X}^{(k)}_{N}={\bf M}^{(k)}_{N}\odot{\bf X}$. Note that knowledge of the intent matrix $\bf I$} is assumed at the transmitter. Our proposed framework takes a mixed approach by using a conventional lossy compression scheme for the intended portion of the source, and generative AI for the other semantic classes, i.e. the non-intended portion, at the receiver. Hence, the signal retrieved at each user is partially reconstructed and partially synthesized with generative AI based on the multi-user intent. This allows to maintain a high fidelity for the intended classes while significantly reducing the communication overhead by transmitting the other classes over shared wireless resource at a low rate with a generative scheme.
@@ II-A2 Mixed Reconstruction/Synthesis with Semantic Diffusion Models
At the receiver, each user receives two datastreams, one containing its intended classes which is decompressed to reconstruct its intended portion of the source signal, and the other containing the semantic map $\bf M$} which is fed into a pre-trained generative diffusion model for high-fidelity synthesis of its non-intended portion. The signal retrieved at user $k$ is thereby written as ${\bf Y}^{(k)}={\bf Y}^{(k)}_{r}\bigcup{\bf Y}^{(k)}_{s}$, where ${\bf Y}^{(k)}_{r}$ is the reconstructed and ${\bf Y}^{(k)}_{s}$ is the synthesized portion. Any state-of-the-art generative models with semantic guidance can be used to locally synthesize the non-intended signal portions ${\bf Y}^{(k)}_{s}$ at the users. Here we pre-train GDMs with classifier-free guidance for high quality/fidelity image synthesis from the semantic map at each user. Note that we adopt diffusion models in the proposed framework as they enable synthesizing photorealistic outputs with better diversity, realism, and fine-grained details, thereby significantly improving users’ quality of experience, e.g. in comparison with other classes of generative models, e.g. GANs, VAEs, etc. We provide comparisons with GAN-based synthesis of the non-intended classes later in simulation results. Refer to [31, 32, 33, 34, 35] for details on diffusion models and their various recent applications in wireless communications.
@@
where ${\bf N}\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\sigma_{t}$ is a trainable parameter controlling the noise scale at each step. For more information on diffusion models and semantic guidance refer to [37].
@@ II-B1 Multicasting the Semantic Map over Shared Wireless Resources
The semantic map is compressed by ${\bf b}_{0}=\mathcal{C}({\bf M};r_{0})$, where $\mathcal{C}(.;r_{0})$ denotes the compression operator, $r_{0}$ is an adaptable compression rate in bit per pixel (bpp), and ${\bf b}_{0}$ is the resulting datastream which is then multicasted to all the users over shared wireless resources. [^1] Various compression techniques, e.g. JPEG[38], BPG [39], NTC [40], etc., can be used based on the type of the source signal to be transmitted and design requirements. The compression rate $r_{0}$ controls the resulting datasize of the semantic map, i.e. ${\left|{{{\bf{b}}_{0}}}\right|}=r_{0}\times{\left|{\bf X}\right|}$, and is adapted based on the wireless channel conditions and the received signal-to-noise ratios (SNRs) of the users. The resulting datastream is then encoded and modulated over the channel, with $\mathcal{CM}(.)$ representing the modulation and coding operator. Denoting the transmit power of the semantic map by $p_{0}$, the resulting data rate of the semantic map at user $k$ is given by
@@ II-B2 Orthogonal Transmission of the User Intended Classes
The reconstruction and synthesis distortion/perception requirements are measured by $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ metrics, respectively. These can be any distortion or perception metrics, e.g. MS-SSIM [41], LPIPS [42], FID [43], etc. According to the rate-distortion-perception theorem [12], $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ are decreasing functions of the compression rates, i.e. $r_{l},r_{0}$. Minimization of the sum users latency with the intent matrix is thereby formulated as
@@
It is also straightforward to show that optimization (8) is convex, as the reconstruction/synthesis distortion/perception curves $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ are monotonically non-increasing according to the rate-distortion-perception theorem [11, 12]. This optimization is a constrained non-linear problem, which we solve using sequential quadratic programming (SQP) [44]. Let us define ${\bf z}=[p_{0},\cdots,p_{L},r_{0},\cdots,r_{L},z_{1},\cdots,z_{K}]^{\top}$, and
@@
4Initialize Set: $i=0$, Randomize: $\mathbf{z}_{0}$$\{\alpha_{k}\}$, $\{\beta_{kl}\}$, $\{\mu_{l}\}$, $\delta$, $\sigma$, Calculate: $\mathcal{E}^{(r)}_{l}=\min\left\{\varepsilon_{kl}^{(r)}\big{|}[{\bf I}]_{kl}\neq 0,k\in[K]\right\},~\forall l\in[L]$, $\mathcal{E}^{(s)}=\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}$.
5
6while $i<\mathrm{Iter}_{\mathrm{max}}$ and $\|\nabla f(\mathbf{z}_{i})\|>\mathrm{Tol}$ do
@@ III-A Reconstruction/Synthesis Rate Distortion/Perception Curves
In this work, we adopt the MS-SSIM and LPIPS as the distortion and perception metrics to evaluate quality of the reconstructed and synthesized signals at the users, respectively. MS-SSIM (Multi-Scale Structural Similarity Index) is based on the pixel-level differences [41], while LPIPS (Learned Perceptual Image Patch Similarity) is computed by taking the L2 norm of the differences between feature maps from different layers of a pre-trained VGG network, normalized and weighted by layer importance [42]. LPIPS is an effective metric for assessing the perceptual quality of images synthesized by generative models, super-resolution, style transfer, etc. and is known to offer a human-like evaluation.
@@ III-B The Simulation Setup
We consider a generative semantic multicasting setup, where a transmitter equipped with a camera multicasts the street scene images to multiple users based on the multi-user intent. The street scene images are taken randomly from the Cityscapes dataset. Cityscapes is a benchmark for semantic understanding of urban street scenes [45], which includes 8 different semantic categories, namely, construction, object, nature, sky, human, vehicle, flat, and void, and a total of 35 semantic object classes. Each category may include several object classes, e.g., the vehicle category covers “Bicycle”, “Motorcycle”, “Train”, “Trailer”, “Caravan”, and “Car” object classes. Error-free segmentation of the scene images at transmitter is assumed. For now, we assume each user is interested in a random class of the 35, with no two users interested in the same class, i.e. non-overlapped intents. Later in Subsection IV.H, we investigate effects of overlapped intents as well. The wireless transmission parameters are listed in TABLE II. All results are averaged over 6000 Monte Carlo simulations.
@@ III-C Comparison with Non-Generative and Full Generative Intent-Unaware Multicasting Benchmarks
2) Full Generative intent-unaware Multicasting (FGM) [37]: In this benchmark, the transmitter compresses the semantic map at rate $r_{FGM}=\Phi_{s}^{-1}\left(\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}\right)$ and multicasts it to all users over shared resources, where each user then carries out SDM according to [37], to locally synthesize a signal utilizing its received semantic map.

### v3

@@ I Introduction
Recently, the Generative AI (GenAI) models have shown great success in developing efficient low bitrate Generative Semantic Communication (Gen Semcom) systems [Li2024generative, xia2023generative, AR, Ece]. Generative models are capable of learning the general distribution of natural signals during training and synthesizing new samples with high perceptual quality at inference time. The generation process can be guided by multi-modal prompts and conditioning signals to produce high quality outputs with a desired semantic content. The emerging Generative Foundation Models and Multimodal Large Language Models (MLLMs), e.g. Sora, Lumiere, and DALL.E, etc., provide ample opportunities to develop efficient generative Semcom frameworks [TokCom, xu2024semantic, Foundation1, Large, LightweightLAI]. As these models possess a vast general knowledge captured via intensive pre-training on huge amount of data, they alleviate the need for a shared Knowledge Base (KB) between the semantic transmitter and receiver, thereby reducing the corresponding knowledge sharing overheads. This vast general knowledge also makes the generative Semcom framework applicable to various datasets and tasks thereby achieving universality. The GenAI models are trained to maximize the perceptual quality of the synthesized signal, and the theoretical limits of generative Semcom are governed by the rate-distortion-perception theory [RDP, RDP2]. This theory explores the threefold trade-off between rate, distortion, and the perceptual quality of the synthesized signal. While the rate-distortion-perception function is analytically derived only for a few source distributions, it is generally estimated empirically for natural signals, e.g. images, audio/video, point cloud, etc., by training deep source encoders with a perceptual loss function, e.g. Wasserstein distance, on large datasets. In generative Semcom, the transmitter extracts the intended semantics, e.g. in form of textual prompts [Prompt1, Prompt2, Genvideosem], compressed embeddings [Giordano2024generative, yilmaz2024high], semantic map [LightDiff, WU2024519], edge map [Li2024generative, xu2024semantic], etc., which are then transmitted over the channel. The receiver uses these semantics to guide a generative model, synthesizing a signal that is semantically consistent and highly realistic. Despite the above works, multiuser generative semantic communications, specifically in the multicasting setup, and its scalability to increased number of users, considering the communication/computation latency remains less studied.
[R1] [R2] [R3] [R4] This Work
@@
In this paper, we develop an scalable latency-aware generative semantic multicasting framework with adaptive intent-based resource allocation utilizing pre-trained diffusion models. We claim three key benefits for the proposed framework: Firstly, generative multicasting in our Semcom framework improves utilization of the wireless and network resources compared with non-generative and intent-unaware multicasting benchmarks. This also reduces the total latency while maintaining the reconstruction/synthesis distortion/perception quality. Secondly, the adoption of pre-trained models allows a source-channel separation-based Semcom architecture, thereby alleviating the need for end-to-end joint training of the transmitter and receiver, which is required in many existing Semcom frameworks [SemCom1, Semcom2, pointcloudsem, SemCom4]. Such a separation-based architecture offers improved adaptability to varying channel conditions, improved scalability to increase the number of users, and better compatibility with the existing design of wireless communication networks. Finally, diffusion models enable synthesizing photorealistic outputs with better diversity, realism, and fine-grained details. TABLE I provides an overview of the major differences between this work and the key existing works on semantic multicasting/broadcasting [R1, R2, R3, R4]. To the best of our knowledge, this work is the first of its kind to design joint communication/computation latency-aware intent-based generative semantic multicasting with pre-trained diffusion models, based on source-channel coding separation to achieve scalable and adaptable semantic multicasting compatible with the existing wireless networks, and derive the corresponding rate-distortion/perception tradeoffs. The contributions of this work are thereby three-fold:
@@ II-A1 Multi-class Semantic Decomposition with DDRNets
At the transmitter, a semantic segmentation model is used to generate a semantic map $\bf M$} from the source signal as depicted in Fig. 1 for sample image signals. To do this, any existing semantic segmentation model, e.g. [Seg2, Seg1, SAM1], can be utilized, and the choice of the appropriate model depends on the required accuracy and affordable computational complexity. To minimize the end-to-end latency, we use a deep dual-resolution network (DDRNet)[Seg1]-based architecture. DDRNets are efficient architectures that were proposed to achieve real-time performance for semantic segmentation of the street scenes. DDRNets are composed of two deep convolutional branches, and a deep information extractor called Deep aggregation pyramid pooling module (DAPPM) to enlarge effective receptive fields and fuse multi-scale context based on low-resolution feature maps as shown in Fig. 2. RB and RBB represent sequential and single residual bottleneck blocks. Information is extracted from low-resolution feature maps e.g., $1/8$ denotes high-resolution branch create feature maps whose resolution is 1/8 of the input image resolution. The black solid lines demonstrate the data processing path (including up-sampling and down-sampling) and black dashed lines denote information paths without data processing. The architecture is trained with the cross entropy loss to classify each image pixel into a semantic class, thereby generating the semantic map $\bf M$}. The semantic map is then converted to multiple one-hot binary masks ${\bf M}_{l}$. With this representation, we have ${\bf X}_{l}={\bf M}_{l}\odot{\bf X}$. The intended mask of user $k$ is then ${\bf M}^{(k)}_{I}=\{\bigcup_{l}{\bf M}_{l}|[{\bf I}]_{kl}=1\}$ and we have ${\bf X}^{(k)}_{I}={\bf M}^{(k)}_{I}\odot{\bf X}$. Similarly, the non-intended mask of user $k$ is then ${\bf M}^{(k)}_{N}=\{\bigcup_{l}{\bf M}_{l}|[{\bf I}]_{kl}=0\}$ and we have ${\bf X}^{(k)}_{N}={\bf M}^{(k)}_{N}\odot{\bf X}$. Note that knowledge of the intent matrix $\bf I$} is assumed at the transmitter. Our proposed framework takes a mixed approach by using a conventional lossy compression scheme for the intended portion of the source, and generative AI for the other semantic classes, i.e. the non-intended portion, at the receiver. Hence, the signal retrieved at each user is partially reconstructed and partially synthesized with generative AI based on the multi-user intent. This allows to maintain a high fidelity for the intended classes while significantly reducing the communication overhead by transmitting the other classes over shared wireless resource at a low rate with a generative scheme.
@@ II-A2 Mixed Reconstruction/Synthesis with Semantic Diffusion Models
At the receiver, each user receives two datastreams, one containing its intended classes which is decompressed to reconstruct its intended portion of the source signal, and the other containing the semantic map $\bf M$} which is fed into a pre-trained generative diffusion model for high-fidelity synthesis of its non-intended portion. The signal retrieved at user $k$ is thereby written as ${\bf Y}^{(k)}={\bf Y}^{(k)}_{r}\bigcup{\bf Y}^{(k)}_{s}$, where ${\bf Y}^{(k)}_{r}$ is the reconstructed and ${\bf Y}^{(k)}_{s}$ is the synthesized portion. Any state-of-the-art generative models with semantic guidance can be used to locally synthesize the non-intended signal portions ${\bf Y}^{(k)}_{s}$ at the users. Here we pre-train GDMs with classifier-free guidance for high quality/fidelity image synthesis from the semantic map at each user. Note that we adopt diffusion models in the proposed framework as they enable synthesizing photorealistic outputs with better diversity, realism, and fine-grained details, thereby significantly improving users’ quality of experience, e.g. in comparison with other classes of generative models, e.g. GANs, VAEs, etc. We provide comparisons with GAN-based synthesis of the non-intended classes later in simulation results. Refer to [CDDM, Dawn, Reinforcement, Matti, SurveyDiff] for details on diffusion models and their various recent applications in wireless communications.
@@
where ${\bf N}\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\sigma_{t}$ is a trainable parameter controlling the noise scale at each step. For more information on diffusion models and semantic guidance refer to [SDM2025].
@@ II-B1 Multicasting the Semantic Map over Shared Wireless Resources
The semantic map is compressed by ${\bf b}_{0}=\mathcal{C}({\bf M};r_{0})$, where $\mathcal{C}(.;r_{0})$ denotes the compression operator, $r_{0}$ is an adaptable compression rate in bit per pixel (bpp), and ${\bf b}_{0}$ is the resulting datastream which is then multicasted to all the users over shared wireless resources. [^1] Various compression techniques, e.g. JPEG[JPEG], BPG [BPG], NTC [balle2020nonlinear], etc., can be used based on the type of the source signal to be transmitted and design requirements. The compression rate $r_{0}$ controls the resulting datasize of the semantic map, i.e. ${\left|{{{\bf{b}}_{0}}}\right|}=r_{0}\times{\left|{\bf X}\right|}$, and is adapted based on the wireless channel conditions and the received signal-to-noise ratios (SNRs) of the users. The resulting datastream is then encoded and modulated over the channel, with $\mathcal{CM}(.)$ representing the modulation and coding operator. Denoting the transmit power of the semantic map by $p_{0}$, the resulting data rate of the semantic map at user $k$ is given by
@@ II-B2 Orthogonal Transmission of the User Intended Classes
The reconstruction and synthesis distortion/perception requirements are measured by $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ metrics, respectively. These can be any distortion or perception metrics, e.g. MS-SSIM [SSIM], LPIPS [LPIPS], FID [FID], etc. According to the rate-distortion-perception theorem [RDP2], $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ are decreasing functions of the compression rates, i.e. $r_{l},r_{0}$. Minimization of the sum users latency with the intent matrix is thereby formulated as
@@
It is also straightforward to show that optimization (8) is convex, as the reconstruction/synthesis distortion/perception curves $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ are monotonically non-increasing according to the rate-distortion-perception theorem [RDP, RDP2]. This optimization is a constrained non-linear problem, which we solve using sequential quadratic programming (SQP) [NoceWrig06]. Let us define ${\bf z}=[p_{0},\cdots,p_{L},r_{0},\cdots,r_{L},z_{1},\cdots,z_{K}]^{\top}$, and
@@
41ex
5Initialize Set: $i=0$, Randomize: $\mathbf{z}_{0}$$\{\alpha_{k}\}$, $\{\beta_{kl}\}$, $\{\mu_{l}\}$, $\delta$, $\sigma$, Calculate: $\mathcal{E}^{(r)}_{l}=\min\left\{\varepsilon_{kl}^{(r)}\big|[{\bf I}]_{kl}\neq 0,k\in[K]\right\},~\forall l\in[L]$, $\mathcal{E}^{(s)}=\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}$.
6
7while $i<\mathrm{Iter}_{\mathrm{max}}$ and $\|\nabla f(\mathbf{z}_{i})\|>\mathrm{Tol}$ do
@@ III-A Reconstruction/Synthesis Rate Distortion/Perception Curves
In this work, we adopt the MS-SSIM and LPIPS as the distortion and perception metrics to evaluate quality of the reconstructed and synthesized signals at the users, respectively. MS-SSIM (Multi-Scale Structural Similarity Index) is based on the pixel-level differences [SSIM], while LPIPS (Learned Perceptual Image Patch Similarity) is computed by taking the L2 norm of the differences between feature maps from different layers of a pre-trained VGG network, normalized and weighted by layer importance [LPIPS]. LPIPS is an effective metric for assessing the perceptual quality of images synthesized by generative models, super-resolution, style transfer, etc. and is known to offer a human-like evaluation.
@@ III-B The Simulation Setup
We consider a generative semantic multicasting setup, where a transmitter equipped with a camera multicasts the street scene images to multiple users based on the multi-user intent. The street scene images are taken randomly from the Cityscapes dataset. Cityscapes is a benchmark for semantic understanding of urban street scenes [Cordts2016Cityscapes], which includes 8 different semantic categories, namely, construction, object, nature, sky, human, vehicle, flat, and void, and a total of 35 semantic object classes. Each category may include several object classes, e.g., the vehicle category covers “Bicycle”, “Motorcycle”, “Train”, “Trailer”, “Caravan”, and “Car” object classes. Error-free segmentation of the scene images at transmitter is assumed. For now, we assume each user is interested in a random class of the 35, with no two users interested in the same class, i.e. non-overlapped intents. Later in Subsection IV.H, we investigate effects of overlapped intents as well. The wireless transmission parameters are listed in TABLE II. All results are averaged over 6000 Monte Carlo simulations.
@@ III-C Comparison with Non-Generative and Full Generative Intent-Unaware Multicasting Benchmarks
2) Full Generative intent-unaware Multicasting (FGM) [SDM2025]: In this benchmark, the transmitter compresses the semantic map at rate $r_{FGM}=\Phi_{s}^{-1}\left(\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}\right)$ and multicasts it to all users over shared resources, where each user then carries out SDM according to [SDM2025], to locally synthesize a signal utilizing its received semantic map.